# CSI-Bench: A Large-Scale In-the-Wild Dataset for Multi-task WiFi Sensing

### v1

@@ 2.1 WiFi Sensing
Compared to vision-, audio-, or wearable-based systems, WiFi sensing offers a scalable, privacy-preserving, and non-intrusive alternative or complementary solution for continuous monitoring in smart environments and healthcare applications Wi-Fi2020 ; electronics12234858 ; SOTO202299 . WiFi sensing has demonstrated substantial potential in tasks such as activity recognition bing2021that ; Meneghello2023activity , gesture detection sai2023gesture ; yang2019gesture , indoor localization wang2017localization ; wang2024localization , and vital sign monitoring wang2024vital ; gao2020vital . However, most existing studies rely on data collected in constrained settings, which limits generalization to diverse users, hardware platforms, and real-world deployment scenarios.
@@ 3.1 Overview
To support robust and generalizable WiFi sensing research, we build a diverse collection of datasets captured in real-world environments using commercial WiFi devices. CSI-Bench spans over 460 hours of CSI recordings across 35 unique users, 26 environments, and 16 device types, covering both routers and edge devices operating under varied network conditions. Data is collected in homes, offices, and public indoor areas with minimal control over ambient interference or user behavior. Each dataset is designed to support one or more sensing tasks, including fall detection (Fall), breathing monitoring (Breath), localization (Loc.), human activity recognition (HAR), user identification (UID), and proximity estimation (Prox.). Representative CSI samples illustrating task-specific signal patterns are visualized in Figure 2. The following section details the hardware, environments, and collection protocols used to capture the datasets.
@@ 3.2 Devices and Hardware Setup
Hardware. To emulate the heterogeneity of real-world WiFi sensing deployments, we select a diverse set of WiFi routers and edge IoT devices commonly found in residential and commercial environments, with chipset models including Qualcomm, NXP, Broadcom, and Espressif Qualcomm ; NXP ; Broadcom ; Espressif . All devices collectively support IEEE 802.11n/ac/ax standards, with MIMO configurations ranging from 1×1 to 2×2 and 1×4, and channel bandwidths of 20, 40, and 80 MHz. The detailed specifications of the edge IoT devices are provided in Appendix. A.
@@ 3.3 Continuous Data Recording
We also develop a lightweight user annotation tool integrated into Google Spreadsheet GoogleSheets , allowing users to optionally log daily activities—such as waking up, sleeping, leaving or returning home, room occupancy, or inactivity—by tapping buttons that record local timestamps. Our system queries and retrieves CSI files matching these events, concatenates the relevant segments, and refines alignment using embedded packet-level timestamps, resulting in precisely labeled CSI data segments. We collect CSI of motion from non-human sources like pets and cleaning robot when users are not home. When possible, time-aligned external information is collected through camera recordings and local sensor logs to annotate non-human motions or highlight environmental changes.
@@ 4.2 CSI Preprocessing Pipeline
Data segmentation. To facilitate task-specific model training, we segment the collected CSI data into fixed-duration samples. For tasks including Fall Detection, Localization, Motion Source Recognition, and the Multi-Task dataset, we segment CSI data into 5-second intervals. For the Breathing Detection dataset, considering the slower temporal variations inherent to respiration signals, we segment the CSI data into 10-second intervals.
Amplitude normalization. The automatic gain controller on commercial WiFi devices affects the reported CSI amplitude. To mitigate the effects of varying signal strengths, we normalize the power response of each subcarrier across the entire frequency band as follows $\hat{H}(f_{k},t)=\frac{|H(f_{k},t)|^{2}}{\sum^{N_{s}}_{k=1}|H(f_{k},t)|^{2}},\ \mathrm{for}\ \mathrm{all}\ k,$ where $N_{s}$ is the number of subcarriers, and $H(f_{k},t)$ is the original reported CSI on the $k$-th subcarrier.
@@ 5.3 Baseline Models
Acc F1 Acc F1 $\Delta$Acc $\Delta$F1
Human activity Recognition 75.40 ±0.93 75.49 ±0.73 87.79 ±0.00 86.47 ±0.00 +12.39 +10.98
User Identification 99.51 ±0.32 99.51 ±0.32 99.83 ±0.00 100.00 ±0.00 +0.32 +0.49
Proximity Recognition 77.52 ±3.13 77.35 ±3.24 87.85 ±0.00 88.67 ±0.00 +10.33 +11.32
5.4 Results
@@
In addition to strong performance, our multi-task framework significantly reduces model complexity and training cost. By consolidating three single-task Transformers into a single backbone with task-specific adapters, we reduce the total parameter count by over 60%. This compression is achieved without degrading task performance. Moreover, because all tasks are trained jointly in a single pass, the wall-clock training time is reduced by nearly 3×3$\times$3 × compared to training separate models for each task. These gains in model size and training efficiency make our approach especially suitable for deployment on resource-constrained edge devices, where memory and compute budgets are limited.
@@
While models perform well under in-distribution settings, we observe significant performance degradation under domain shifts. OOD evaluation across user, environment, and device axes—summarized in Appendix C.2—reveals notable challenges in generalization, particularly in cross-device settings. This highlights a key motivation for developing robust and adaptive models in future work.
@@ 7 Conclusion
CSI-Bench enables privacy-preserving, contactless health monitoring using commercial WiFi, supporting applications like fall detection and breathing monitoring in homes and care facilities. By providing a large-scale, real-world dataset with diverse users, devices, and environments, it fosters inclusive, reproducible research and robust model development. Its high-dimensional CSI data presents unique challenges and opportunities for advancing ML methods in time-series analysis, representation learning, and domain generalization. All data were collected with consent, anonymized, and ethically managed. CSI-Bench serves both as a foundation for practical healthcare solutions and as a benchmark for high-dimensional, human-centered AI systems.
@@ A.3 Devices and Hardware Diversity
To ensure broad coverage of real-world IoT infrastructure, we select 16 types of commercial WiFi-enabled devices operating across both 2.4 GHz and 5 GHz bands, with bandwidths of 20, 40, and 80 MHz. Devices span major chipset vendors such as Qualcomm, Broadcom, Espressif, and NXP. Prior to data collection, each device is evaluated using our in-house CSI verification tool to assess signal consistency, sampling stability, and amplitude dynamics. Devices that pass quality thresholds are used in deployment. Figure 4 presents the CSI quality scores across candidate devices; Table 6 lists their specifications.
@@ A.4 Task-Specific Dataset Statistics
Device WiFi Chipset Vendor Model Antenna Bandwidth Band
AmazonPlug MediaTek MT7697N 1x1 20MHz 2.4G
GoveePlug Espressif ESP8266/ESP8285 1x1 20MHz 2.4G
WyzePlug Espressif ESP8266/ESP8285 1x1 20MHz 2.4G
EightreePlug Espressif ESP8266/ESP8285 1x1 20MHz 2.4G
EchoPlus MediaTek MT8516 1x1 20/40/80MHz 2.4G & 5G
GoogleNest Qualcomm IPQ4019 1x1 20/40MHz 2.4G & 5G
AppleHomePod - - 1x1 20/40MHz 2.4G & 5G
EchoSpot MediaTek MT6625L 1x1 20/40MHz 2.4G & 5G
EchoShow 8 MediaTek MT8183 1x1 20/40/80MHz 2.4G & 5G
Echodot 2 gen MediaTek MT6625LN 1x1 20/40MHz 2.4G & 5G
Echodot 3 gen MediaTek MT7658CSN 1x1 20/40/80MHz 2.4G & 5G
Hex Home Qualcomm - 1x2 20/40MHz 5G
@@
ESP32 Espressif S3 1x1 20/40MHz 2.4G
GoogleNestHub Broadcom BCM4345 1x1 20/40/80MHz 2.4G & 5G
Lyra Qualcomm - 2x2 20/40/80MHz 2.4G & 5G
@@
E03 Apartment 1077 2B2B 1
E04 Apartment 790 1B1B 1
E05 Apartment 714 1B1B 1
@@ B.3 Training Protocol
All models are trained with the AdamW optimizer, a batch size of 128, and initial learning rate of $31e-$ 3. We apply cosine learning rate decay with 5 warm-up epochs and weight decay of $51e-$ 5. Training lasts up to 100 epochs, with early stopping based on validation loss (patience = 15). We use categorical cross-entropy as the loss function. The hyperparameter are tuned based on models’ accuracy on validation dataset. Data is loaded from HDF5 using standardized splits as discussed in Section 5.2 and label mappings. Our experiments utilize NVIDIA GeForce RTX 4090 GPUs and AWS Sagemaker involved training with three random seeds across all datasets. For training in AWS Sagemaker, we use ml.g5.g5.12xlarge, which includes 4 NVIDIA A10G Tensor Core GPUs. The training time for tasks ranges from 0.5 hour to 13 hours.

### v2

@@ 2.1 WiFi Sensing
Compared to vision-, audio-, or wearable-based systems, WiFi sensing offers a scalable, privacy-preserving, and non-intrusive alternative or complementary solution for continuous monitoring in smart environments and healthcare applications Wi-Fi2020; electronics12234858; SOTO202299; SrcSense. WiFi sensing has demonstrated substantial potential in tasks such as activity recognition bing2021that; Meneghello2023activity, gesture detection sai2023gesture; yang2019gesture, indoor tracking and localization wang2017localization; wang2024localization; EZMap; floorplan, fall detection FallAware, proximity detection proximity, and vital sign monitoring wang2024vital; gao2020vital. However, most existing studies rely on data collected in constrained settings, which limits generalization to diverse users, hardware platforms, and real-world deployment scenarios.
@@ 3.1 Overview
To support robust and generalizable WiFi sensing research, we build a diverse collection of datasets captured in real-world environments using commercial WiFi devices. CSI-Bench spans over 461 hours of CSI recordings across 35 unique users, 26 environments, and 16 device types, covering both routers and edge devices operating under varied network conditions. Data is collected in homes, offices, and public indoor areas with minimal control over ambient interference or user behavior. Each dataset is designed to support one or more sensing tasks, including fall detection (Fall), breathing monitoring (Breath), localization (Loc.), human activity recognition (HAR), user identification (UID), and proximity estimation (Prox.). Representative CSI samples illustrating task-specific signal patterns are visualized in Figure 2. The following section details the hardware, environments, and collection protocols used to capture the datasets.
@@ 3.2 Devices and Hardware Setup
Hardware. To emulate the heterogeneity of real-world WiFi sensing deployments, we employ a diverse set of commercial WiFi routers and edge IoT devices commonly found in residential and commercial environments. The selected devices span five major chipset vendors, including Qualcomm, MediaTek, Broadcom, Espressif, and NXP Qualcomm; NXP; Broadcom; Espressif, and cover a broad spectrum of hardware configurations, including 1×1 to 2×2 MIMO and 1×4 antenna setups. All devices support IEEE 802.11 n/ac/ax standards, operating across both 2.4 GHz and 5 GHz bands with channel bandwidths of 20, 40, and 80 MHz. These heterogeneous devices are intentionally chosen to reflect the real IoT ecosystem deployed in homes, offices, and small businesses, where heterogeneous devices with varying wireless capabilities coexist in complex indoor environments. Their detailed specifications, including chipset model, antenna configuration, bandwidth, frequency band, and empirically measured average RSSI, are listed in Appendix A (Table 7).
@@ 3.3 Continuous Data Recording
We also develop a lightweight user annotation tool integrated into Google Spreadsheet GoogleSheets, allowing users to optionally log daily activities—such as waking up, sleeping, leaving or returning home, room occupancy, or inactivity—by tapping buttons that record local timestamps. This design minimizes user effort while ensuring accurate temporal alignment between activity logs and CSI data. An illustration of the annotation tool is provided in Appendix A.10. Our system queries and retrieves CSI files matching these events, concatenates the relevant segments, and refines alignment using embedded packet-level timestamps, resulting in precisely labeled CSI data segments. We collect CSI of motion from non-human sources like pets and cleaning robot when users are not home. When possible, time-aligned external information is collected through camera recordings and local sensor logs to annotate non-human motions or highlight environmental changes.
@@ 4.2 CSI Preprocessing Pipeline
Data segmentation. To facilitate supervised model training, the collected CSI data is segmented into fixed-duration, non-overlapping samples. For tasks including Fall Detection, Localization, Motion Source Recognition, and the Multi-Task dataset, we segment CSI data into 5-second intervals. For the Breathing Detection dataset, considering the slower temporal variations inherent to respiration signals, we segment the CSI data into 10-second intervals. We also provide the unsegmented CSI recordings of part of our dataset to support
Amplitude normalization. To mitigate the effects of varying signal strengths, we normalize each CSI sample across all subcarriers and time indices by removing the mean and scaling by the standard deviation. This ensures consistent scaling across samples while preserving the relative temporal–spectral dynamics within each sample. The normalized CSI is computed as $\hat{H}(f_{k},t)=\frac{H(f_{k},t)-\mu_{H}}{\sigma_{H}+\epsilon}$ where $\mu_{H}$ and $\sigma_{H}$ denote the mean and standard deviation of $H(f_{k},t)$ over the entire CSI matrix, and $\epsilon$ is a small constant added for numerical stability.
@@ 5.3 Baseline Models
Acc F1 Acc F1 $\DeltaAcc Δ\DeltaF$1
Human Activity Recognition 75.40 ±0.93 75.49 ±0.73 88.06 ±0.76 86.00 ±2.05 +12.66 +10.51
User Identification 99.51 ±0.32 99.51 ±0.32 99.55 ±0.06 99.70 ±0.27 +0.04 +0.19
Proximity Recognition 77.52 ±3.13 77.35 ±3.24 86.41 ±1.97 87.09 ±1.46 +8.89 +9.74
5.4 Baseline Evaluation
@@
In addition to strong performance, our multi-task framework significantly reduces model complexity and training cost. By consolidating three single-task Transformers into a single backbone with task-specific adapters, we reduce the total parameter count by over 60%. This compression is achieved without degrading task performance. Moreover, because all tasks are trained jointly in a single pass, the wall-clock training time is reduced by nearly $3\times$ compared to training separate models for each task. These gains in model size and training efficiency make our approach especially suitable for deployment on resource-constrained edge devices, where memory and compute budgets are limited.
@@
5.5 Evaluation on OOD Splits
We evaluate the Transformer-based multi-task model under three OOD conditions: cross-device, cross-environment, and cross-user (Table 5). Compared with the in-distribution results in Table 4, all OOD accuracies and F1-scores drop substantially, indicating that models trained on seen domains fail to generalize effectively to unseen users, environments, or hardware. This degradation reflects WiFi CSI’s strong sensitivity to device, environmental, and user variations, revealing a significant generalization gap between in-distribution and OOD domains. The results highlight the need for domain-adaptive, calibration-free learning frameworks to improve real-world robustness in WiFi sensing. More detailed results and analysis can be found in Appendix C.2.
Table 5: Cross-domain performance of Transformer model on three tasks. Accuracy (Acc) and F1-score are reported as mean ± std (%) over three runs.
Task Cross-Device Cross-Env Cross-User
Acc F1 Acc F1 Acc F1
Human Activity Recognition 61.82 ±0.95 57.80 ±0.78 54.92 ±0.98 47.17 ±1.12 54.72 ±0.84 46.67 ±1.00
Human Identification 59.94 ±0.77 59.81 ±0.96 / / / /
Proximity Recognition 30.68 ±3.11 28.76 ±3.51 29.67 ±1.63 27.12 ±1.79 30.26 ±1.94 25.97 ±2.26
@@ 7 Conclusion
CSI-Bench establishes a foundational step toward scalable, privacy-preserving, and contactless healthcare enabled by commodity WiFi infrastructure. By leveraging signals already emitted in everyday environments, it supports continuous health and behavioral monitoring without requiring wearables or cameras, thereby reducing barriers to adoption and ensuring dignity, inclusivity, and accessibility in care.
The benchmark supports multiple healthcare-relevant sensing capabilities: fall detection and breathing monitoring for safety and wellness tracking; activity recognition and localization for rehabilitation and behavioral analysis; user identification and proximity estimation for personalized assistance and social-interaction assessment; and motion source recognition to reduce false alarms in automated home-care systems. Collectively, these tasks lay the groundwork for comprehensive, unobtrusive smart health monitoring in homes, hospitals, and assisted living facilities.
Beyond practical healthcare applications, CSI-Bench contributes to the broader machine learning and sensing communities by providing a large-scale, in-the-wild dataset that captures the true complexity of human environments. Its diversity in users, hardware, and contexts fosters reproducible research and robust model development for real-world deployment. The high-dimensional CSI signals pose unique challenges in time-series modeling, representation learning, and domain generalization, motivating advances in trustworthy and adaptive AI systems that extend beyond healthcare.
All data are collected with consent, anonymized, and ethically managed. CSI-Bench serves both as a foundation for practical healthcare solutions and as a benchmark for high-dimensional, human-centered AI systems.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Please see Section 6.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: The paper does not include theoretical results.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully discLoSe all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data is provided or not)?
  15. Answer: [Yes]
  16. Justification: The experiments are reproducible. The code, dataset and detailed instructions are provided.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: Please find the links to our code and dataset attached in the abstract. We have detailed instructions on how to use our data and benchmark code included.
  21. Guidelines:
  22. 6.
    Experimental Setting/Details
  23. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  24. Answer: [Yes]
  25. Justification: The training details are discussed in Section  B.3.
  26. 7.
    Experiment Statistical Significance
  27. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  28. Answer: [Yes]
  29. Justification: Results in the paper are reported with error bars representing the standard deviation across three random seeds, please see Table 3 and Table 4 in Section 5.3. The primary sources of variability include random initialization, data shuffling, and train/validation/test splits. These are consistently applied to all baseline and proposed models.
  30. 8.
    Experiments Compute Resources
  31. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  32. Answer: [Yes]
  33. Justification: The computing requirements are discussed in Appendix  B.3.
  34. 9.
    Code Of Ethics
  35. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  36. Answer:[Yes]
  37. Justification: Data were passively collected via commercial WiFi devices in participants’ homes, with informed consent and no direct researcher interaction. An internal ethics review addressed privacy and risk, and all data were anonymized. Participants received fair compensation.
  38. 10.
    Broader Impacts
  39. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  40. Answer: [Yes] Please see Section Broader Impact.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: The paper does not involve the release of models or data that carry a high risk for misuse or dual-use concerns. We do not release any generative models or scraped data from public sources, and the released dataset does not pose foreseeable risks that require additional safeguards.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: We have cited the original papers and repositories that produced the code packages and models used in our work. For each asset, we have included the appropriate references in the paper, along with the license information and relevant URLs where applicable. All assets used are open-source and have been used in compliance with their respective licenses
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We introduce a new large-scale dataset and benchmark code suite for WiFi-based sensing tasks. Please find the links under abstract.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [Yes]
  56. Justification: Our dataset includes data from human subjects in their homes, see 3.5, with written informed consent and fair compensation provided. Consent procedures and compensation details are summarized in the main paper, with full instructions and sample screenshots available in the supplementary material for transparency.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were discLoSed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [Yes]
  60. Justification: Our research involved passive data collection in participants’ homes using commercial WiFi devices, with no direct interaction or intervention, see 3.5.
@@ A.3 Devices and Hardware Diversity
To ensure broad coverage of real-world IoT infrastructure, we select 16 types of commercial WiFi-enabled edge devices operating across both 2.4 GHz and 5 GHz bands, with bandwidths of 20, 40, and 80 MHz. Devices span major chipset vendors such as Qualcomm, Broadcom, Espressif, and NXP, covering configurations from low-cost smart plugs to high-performance routers and smart speakers. Prior to data collection, each device is evaluated using our in-house CSI verification tool to assess signal consistency, sampling stability, and amplitude dynamics. Devices that pass quality thresholds are used in deployment. Figure 4 presents the CSI quality scores across candidate devices; Table 7 lists their specifications, together with average RSSI values (in dBm) measured over one-hour static indoor sessions for each device. These RSSI readings characterize real-world signal strength and serve as a practical proxy for transmit-power variability across hardware families.
@@ A.4 Task-Specific Dataset Statistics
Device Chipset Model Antenna Bandwidth Band RSSI
AmazonPlug MediaTek MT7697N 1x1 20MHz 2.4G -49.95 dBm
GoveePlug Espressif ESP8266/ESP8285 1x1 20MHz 2.4G -50.26 dBm
WyzePlug Espressif ESP8266/ESP8285 1x1 20MHz 2.4G -55.57 dBm
EightreePlug Espressif ESP8266/ESP8285 1x1 20MHz 2.4G -54.01 dBm
EchoPlus MediaTek MT8516 1x1 20/40/80MHz 2.4G & 5G -42.24 dBm
GoogleNest Qualcomm IPQ4019 1x1 20/40MHz 2.4G & 5G -49.41 dBm
AppleHomePod 1x1 20/40MHz 2.4G & 5G -52.56 dBm
EchoSpot MediaTek MT6625L 1x1 20/40MHz 2.4G & 5G -51.20 dBm
EchoShow 8 MediaTek MT8183 1x1 20/40/80MHz 2.4G & 5G -38.14 dBm
Echodot 2 gen MediaTek MT6625LN 1x1 20/40MHz 2.4G & 5G -76.83 dBm
Echodot 3 gen MediaTek MT7658CSN 1x1 20/40/80MHz 2.4G & 5G -57.75 dBm
Hex Home Qualcomm 1x2 20/40MHz 5G
@@
ESP32 Espressif S3 1x1 20/40MHz 2.4G -62.86 dBm
GoogleNestHub Broadcom BCM4345 1x1 20/40/80MHz 2.4G & 5G -60.39 dBm
Lyra Qualcomm 2x2 20/40/80MHz 2.4G & 5G
@@
E03 Apartment 1077 $2B$ 1
E04 Apartment 790 $1B$ 1
E05 Apartment 714 $1B$ 1
@@ A.9 Multi-task Dataset
A.10 Annotation Tool
To facilitate user-friendly and accurate labeling during in-the-wild data collection, we developed a lightweight annotation tool based on Google Spreadsheets for accessibility and cross-platform compatibility.
As illustrated in Figure 5, the tool provides a simple interface where users can log activities through “Start/End” button clicks corresponding to predefined motion types (e.g., walking, breathing, jumping, waving hand, running, localization). Each button click automatically records the timestamp, activity label, tester ID, and session duration, ensuring precise temporal alignment with the collected CSI data.
The design emphasizes ease of use and minimal user burden. Participants simply tap the relevant button when starting and finishing an activity—no manual typing or complex input is required. The captured logs include information such as activity type, approximate user–device distance, location context, and timestamps, which are later aligned with the CSI files through automated scripts described in Section 3.3.
During the collection campaign, participants typically logged around three minutes per activity type per day, covering multiple motion categories. While logging was optional, this structured yet flexible protocol ensured sufficient labeled samples for model training while allowing users to behave naturally in their environments.
Figure 5: Screenshot of the Google Spreadsheet–based annotation tool used by participants to record activities and timestamps during data collection.
@@ B.3 Training Protocol
All models are trained with the AdamW optimizer, a batch size of 128, and initial learning rate of 1e31e-3. We apply cosine learning rate decay with 5 warm-up epochs and weight decay of 1e51e-5. Training lasts up to 100 epochs, with early stopping based on validation loss (patience = 15). We use categorical cross-entropy as the loss function. The hyperparameter are tuned based on models’ accuracy on validation dataset. Data is loaded from HDF5 using standardized splits as discussed in Section 5.2 and label mappings. Our experiments utilize NVIDIA GeForce RTX 4090 GPUs and AWS Sagemaker involved training with three random seeds across all datasets. For training in AWS Sagemaker, we use ml.g5.g5.12xlarge, which includes 4 NVIDIA A10G Tensor Core GPUs. The training time for tasks ranges from 0.5 hour to 13 hours.